OpenStudy (anonymous):

Can someone plz help.Simplify the following quotients of complex numbers into the form a + bi. -9 + 8i / 1 + 2i

OpenStudy (nikvist):

\[\frac{-9+8i}{1+2i}=\frac{-9+8i}{1+2i}\frac{1-2i}{1-2i}=\frac{-9+18i+8i+16}{5}=\frac{7+26i}{5}\]

OpenStudy (anonymous):

thank you